# Snail: Secure Single Iteration Localization

### v1

@@ Abstract.
journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
@@ 1. Introduction
Incorporating privacy into visual localization is an active area of research. Consider a lightweight robot equipped with a camera. Privacy preserving localization allows the robot to offload the localization task to more powerful resources without revealing its location, its map, or images taken by its camera. Offloading conserves onboard power and frees resources for other tasks. Privacy is important in this setting because the robot may inadvertently take pictures of sensitive information, or people who do not want their picture taken. Ethical or legal requirements may prevent the images from being shared outside the device domain. Privacy preserving localization also enables settings when the map and camera image are held by different parties, or when both are held by the same party who is trying to prevent data exfiltration by requiring an attacker to compromise multiple systems simultaneously. While technologies have emerged to make privacy preserving localization more efficient, none of the approaches make formal security statements and many of the claimed security properties have been broken (Speciale et al., 2019; Dusmanu et al., 2021; Brachmann and Rother, 2021; Liu et al., 2018).
Previous attempts at secure localization rely on techniques based on obfuscation, like line cloud transformations (Speciale et al., 2019) and adversarial affine subspace embeddings (Dusmanu et al., 2021). Such obfuscation-based techniques do not make formal security statements, but they have low overhead and run efficiently. Consequently, these techniques have been shown to be vulnerable to attacks which can recover the input image (Chelani et al., 2021; Pittaluga and Zhuang, 2023). Recently, differential privacy has been used to address one such attack however privacy is inversely related to the amount of data processed sequentially (Pittaluga and Zhuang, 2023). While this is appropriate for some applications, it is not appropriate for repeated invocation which is common when localizing on sequential camera frames.
@@ 2. Preliminaries
This work builds on the goals of previous works seeking to integrate privacy-enhancing features into localization. Our specific focus is the following two privacy goals: 1) preventing image and map reconstruction and 2) maintaining confidentiality of the pose. These two goals represent the most illustrative standards with which to contrast our methods to the existing state-of-the-art (Speciale et al., 2019; Chelani et al., 2021; Dusmanu et al., 2021). In the remainder of this section we review the techniques used to achieve these goals.
@@ 2.1. Secure Computation
Secure computation describes a cryptographic field that seeks to allow multiple parties to compute a function over secret inputs. Secure multiparty computation (MPC) protocols serve a similar purpose as trusted execution environments (TEEs) but do not require hardware support and are not vulnerable to a class of side channel attacks that affect TEEs (Van Bulck et al., 2017; Lee et al., 2017a, b). When applied in practice, MPC is often collaborative in the sense in that each party has secret inputs to a function, for example a bid in an auction, and the function output, e.g. the winning bid, is learned by all participants. MPC, however, may also be used to offload computation from a weak device to stronger resources such that the weak device with the secret inputs does not participate in the MPC directly, or plays a minor role.
@@
2.2. Visual Localization
Visual localization comprises a series of approaches that infer a device’s location from local features extracted from visual data, traditionally derived from a set of 2D or 3D input images. This process commonly employs structure-based approaches, representing scenes through point clouds and feature matching between retrieved 2D images and 3D point clouds. Point clouds, i.e. maps, may be generated using Structure-from-Motion (SfM) (Yang et al., 2009), constructed simultaneously, or stored from previous mapping of the environment. Such approaches that rely on sharing image features from local 2D images (such as sending them to an external server for analysis) risk leaking private information about the environment to the processing entity, as features are known to be susceptible to image inversion (Weinzaepfel et al., 2011). In practice, several studies demonstrate the recovery of image content from gradient-based feature descriptors, and/or locations, in algorithms such as SIFT (Lowe, 2004) and HOG (Balntas et al., 2017), rendering such methods vulnerable to privacy threats when revealed, as one can reasonably reconstruct the image. A depection is shown in Figure 1.
@@ 2.2. Visual Localization
In the field of computer vision, the PnP problem (Lepetit et al., 2009) consists of estimating the pose (rotation and translation) of a camera with respect to a world coordinate system. The pose is estimated by exploiting a set of 2D feature locations (e.g. points in an image) and their corresponding 3D points, e.g., a 3D map of the environment. If a camera can move freely in 3D space its pose has six degrees of freedom, three rotational and three translational, about the three Cartesian axis. Applications in robotics and AR/VR often use PnP solvers as part of larger processing pipelines including control (Chaumette and Hutchinson, 2006) and mapping (Schonberger and Frahm, 2016). This work focuses only on pose estimation with the PnP problem.
@@ 2.3. Privacy-Preserving Localization
Engelsma et al. (Engelsma et al., 2022) demonstrate a homomorphically encrypted representation search that could be further adapted to existing visual localization algorithms. This approach represents a promising frontier for future privacy-preserving localization algorithms, but is currently in the early stages of application and does not scale to such processes in its current form.
@@ 2.4. Homomorphic Encryption and Localization
The localization algorithms considered are particularly challenging to adapt to execution via homomorphic encryption. Most problematic, is that localization is a high depth computation with significant data dependencies between steps. For example, to perform one iteration of Levenberg-Marquardt localization, over 1000 ciphertext-ciphertext divisions are required with a depth of division operations (sequential divisions) in the low hundreds. Over 30,000 ciphertext-ciphertext multiplications are required with multiplicative depth in the thousands. The high depth of the computation, especially the number of division operations, obviates leveled homomorphic encryption schemes which are much more efficient than the fully homomorphic, bootstrapped schemes including leveled schemes with Residual Number System (RNS) encoding, an optimization for executing high depth computations with leveled schemes. Furthermore, localization has a high degree of dependencies between input data which does not suit the slotting/batching based optimizations to homomorphic encryption schemes popular in BGV and CKKS where multiple plaintexts are encrypted together. Lastly, we later show that floating point precision is required for localization to successfully converge. This obviates the more efficient arithmetic circuit evaluation, requiring encoding the computation as a boolean circuit which is much slower. For these reasons, homomorphic encryption is not further considered and multi-party computation (MPC) techniques are considered instead. While MPC has higher communication complexity than HE, i.e. communication is a function of the size of the computation vs the size of just the inputs, we later show the communication is quite reasonable for MPC-based localization.
@@ 3. Security Model
This definition gives insight into where previous privacy preserving localization attempts have failed. Line cloud transformations of Speciale et. al. cannot be simulated; the lines contained within the view are not known to the simulator and thus the difference between the real and simulated views can be easily distinguished, as practically proven by Chelani et. al. (Speciale et al., 2019; Chelani et al., 2021). Granted, line clouds were only meant to prevent image reconstruction and do not attempt to hide map feature locations M or the output pose $\mathbf{x}$ (see Table 1). However, just considering input feature locations, an adaptation of the above definition to remove the map M and output pose $\mathbf{x}$ still suggests line cloud transformations are not secure. Furthermore, line cloud transformations do not consider the streaming model; authors note their technique is inversely proportional to the amount of data processed in series. A concrete definition of security is needed for future work to build upon, and capturing the streaming model is important to avoid leakage with repeated invocation.
@@ 4. Overview of Privacy Preserving Localization
The second setting considered is where a client holds the features extracted from their image $\mathbf{I}$, but the 3D map $\mathbf{M}$ is held by a third party. The client and third party interact to compute the pose. The client should not learn anything about the map other than what can be inferred from the pose, and the third party should learn (informally) nothing. In the real world, the full map $\mathbf{M}_{\text{full}}$ likely contains more features than the client’s image $\mathbf{I}$. In order to keep $\mathbf{M}_{\text{full}}$ confidential, we assume the existence of an alignment functionality to match the client’s features in $\mathbf{I}$ to the subset of features shared by the map $\mathbf{M}_{full}$, specifically, $\mathbf{M}=align(\mathbf{I},\mathbf{M}_{full})$. We believe the alignment functionality is a reasonably straightforward application of private information retrieval (Chor et al., 1998) and leave concrete instantiation to future work. When alignment is composed with localization, e.g. $x=l(\mathbf{I},\text{align}(\mathbf{I},\mathbf{M}_{\text{full}}))$, the client learns neither the exact map features $\mathbf{M}$, nor map features which did not have a corresponding image feature, e.g. $\mathbf{M}_{\text{full}}\notin\mathbf{M}$.
In both settings considered, the core challenge to instantiating $l$ in a privacy preserving way is in addressing the iterative nature of localization. Iterative algorithms like localization are not friendly to running under MPC because they are not data-oblivious, meaning control flow depends on secret data i.e. the convergence criteria. To run an iterative algorithm under MPC, an upper bound of iterations must be specified a priori, regardless of convergence at runtime, to ensure the rate of convergence is not leaked. If the rate of convergence were leaked, it reveals sensitive information allowing the devices position to be inferred, especially when localization is run repeatedly like on sequential camera frames (as is common in robotics, AR/VR, and autonomous vehicles). The iteration leakage problem has been studied previously in the context of privacy preserving SAT solvers (Luo et al., 2022; Toft, 2009) and there is no direct solution; an upper bound of iterations must be executed to hide the real number of iterations required for convergence. Localization is particularly problematic as it contains an outer gradient descent iterative algorithm, each iteration of which computes the singular value decomposition (SVD) to (pseudo) invert a matrix which is itself an iterative algorithm. Our Single Iteration Localization approach is designed to address the iteration problem, instantiated using the most appropriate localization algorithm and MPC protocol given this context.
@@ 5.1. Localization Algorithm
LM is selected for its wide popularity, real-world usage, and generality as it is the default PnP algorithm in OpenCV (Bradski, 2000) and accepts any number of input point correspondences making it useful in combination with other localization related tasks like RANSAC (Fischler and Bolles, 1981). GN is considered as it requires the same set of linear algebra operations as LM, thus is easy to implement, and also requires fewer matrix multiplications at the cost of inverting a larger matrix. While there are many more algorithmic options available, LM and GN are the most popular and require a narrow variety of operations, i.e., matrix arithmetic 2, point projection 3, Euler finite rotation transformation 3, and a matrix (pseudo) inverse via SVD 4. Furthermore, these algorithms can be directly applied, or easily adjusted, to other computer vision problems like relative pose estimation (Bartoli and Sturm, 2004), homography & image alignment (Hartley and Zisserman, 2004), and Structure-from-Motion (SfM) (Schonberger and Frahm, 2016).
@@ 5.2. MPC Library
5.3. Data Representation
Given promising results from using non-traditional ways to represent data in the field of machine learning, we consider various representations in privacy preserving localization (Gupta et al., 2015; Lesser et al., 2011). Plaintext arithmetic operations on floating point data are accelerated on modern hardware making localization on floating point data fast (around 10 ms for 10k operations) and thus other representations like fixed point are not usually considered. In contrast, when performing floating point operations under the MPC frameworks considered, there are no floating point functional units. Garbled circuit logic gates are evaluated in the same manner no matter how data is represented. For this reason we consider both fixed and floating point data types of various width for privacy preserving localization.
@@ 5.4. Garbled Circuits for Computation Offload
Often, garbled circuit protocols are considered to run between two parties, each of which has secret inputs to the function being computed. In the client-server setting, however, we are considering computation offload where the offload servers have no inputs to the function and are not allowed to learn the output. Achieving this requires some minor logistical modifications to the garbled circuits protocol setup.
Recall garbled circuit protocols have two roles, a generator and an evaluator. The generator creates garbled circuit, which in practice mostly consists of repeatedly evaluating the AES blockcipher. The garbled circuit, i.e. a set of ciphertexts, is then sent to the evaluator who decrypts parts of the received ciphertext, a process that again mostly consists of evaluating AES. The evaluator learns a subset of the ciphertext representing inputs called input wire labels via a primitive named oblivious transfer. The evaluator then uses the input wire labels to decrypt the rest of the garbled circuit.
In the setting where all plaintext inputs and outputs are owned by the client, for whom we additionally want to minimize computation and network usage, the simplest way to get the correct input wire labels from generator to evaluator is to have the generator send all labels to the client, who then forwards the appropriate label per wire onwards to the evaluator. We make a slight optimization from the above approach to reduce the bandwidth requirements of the client. The client chooses a random seed and sends it to the generator who uses a pseudorandom function and the seed to create the wire labels for the garbled circuit. Note that the EMP toolkit framework already uses a seeded PRF to generate the garbled circuit for performance reasons (/dev/random is slow); we have instead allowed the client to choose this seed which is not an issue since the client is the only party who supplies the secret input. After, the generator creates the garbled circuit, they send it to the evaluator excluding the input wire labels. The client, knowing the circuit seed, can generate the input wire labels without interaction and send them to the evaluator. The circuit is evaluated as usual and the output labels are sent back to the client who can decode their plaintext meaning. This saves the client from needing to receive all wire labels from the generator. Instead, the client may generate them directly and send to the evaluator. Since the same seed is used to generate input wire labels and the rest of the garbled circuit, correctness holds. Note in the case of EMP and the half-gates protocol, the client must also know the global circuit delta (Kolesnikov and Schneider, 2008) as part of the generation process. Intuitively, it is okay for the client to learn the seed and delta as the client is the only party who supplies secret inputs and is allowed to learn the output, in the offload setting.

### v2

@@ Abstract.
Multi-Party Computation, Localization, Pose Estimation
journalyear: 2024journalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
@@ 1. Introduction
Incorporating privacy into visual localization is an active area of research. Consider a lightweight robot equipped with a camera. Privacy preserving localization allows the robot to offload the localization task to more powerful resources without revealing its location, its map, or images taken by its camera. Offloading conserves onboard power and frees resources for other tasks. Privacy is important in this setting because the robot may inadvertently take pictures of sensitive information, or people who do not want their picture taken. Ethical or legal requirements may prevent the images from being shared outside the device domain. Privacy preserving localization also enables settings when the map and camera image are held by different parties, or when both are held by the same party who is trying to prevent data exfiltration by requiring an attacker to compromise multiple systems simultaneously. While technologies have emerged to make privacy preserving localization more efficient, none make formal security statements and many of the claimed security properties have been broken (Speciale et al., 2019; Dusmanu et al., 2021; Brachmann and Rother, 2021; Liu et al., 2018).
Previous attempts at secure localization rely on techniques based on obfuscation, like line cloud transformations (Speciale et al., 2019) and adversarial affine subspace embeddings (Dusmanu et al., 2021). Such obfuscation-based techniques do not make formal security statements, but they have low overhead and run efficiently. Consequently, these techniques have been shown to be vulnerable to attacks which can recover the input image (Chelani et al., 2021; Pittaluga and Zhuang, 2023). Recently, differential privacy has been used to address one such attack however privacy is inversely related to the amount of data processed sequentially (Pittaluga and Zhuang, 2023). While this is appropriate for some applications, it is not appropriate for repeated invocation, common when localizing on sequential camera frames.
@@ 2. Preliminaries
This work builds on the goals of previous works seeking to integrate privacy-enhancing features into localization. Our specific focus is the following two privacy goals: 1) preventing image and map reconstruction and 2) maintaining confidentiality of the pose. These two goals represent the most illustrative standards with which to contrast our methods to the existing state-of-the-art (Speciale et al., 2019; Chelani et al., 2021; Dusmanu et al., 2021).
@@ 2.1. Secure Computation
Secure computation describes a cryptographic field that seeks to allow multiple parties to compute a function over secret inputs. Secure multiparty computation (MPC) protocols serve a similar purpose as trusted execution environments (TEEs) but do not require hardware support and are not vulnerable to a class of side channel attacks that affect TEEs (Van Bulck et al., 2017; Lee et al., 2017a, b). When applied in practice, MPC is often collaborative in the sense in that each party has secret inputs to a function, for example a bid in an auction, and the function output, e.g. the winning bid, is learned by all participants. MPC, however, may also be used to offload computation from a weak device to stronger resources such that the weak device with the secret inputs plays a minor role in the protocol.
@@
2.2. Visual Localization
Visual localization comprises a series of approaches that infer a device’s location from local features extracted from visual data, traditionally derived from a set of 2D or 3D input images depicted in Figure 1. This process commonly employs structure-based approaches, representing scenes through point clouds and feature matching between retrieved 2D images and 3D point clouds. Point clouds, i.e. maps, may be generated using Structure-from-Motion (SfM) (Yang et al., 2009), constructed simultaneously, or stored from previous mapping of the environment. Such approaches that rely on sharing image features from local 2D images (such as sending them to an external server for analysis) risk leaking private information about the environment to the processing entity, as features are known to be susceptible to image inversion (Weinzaepfel et al., 2011). In practice, several studies demonstrate the recovery of image content from gradient-based feature descriptors, and/or locations, in algorithms such as SIFT (Lowe, 2004) and HOG (Balntas et al., 2017), rendering such methods vulnerable to privacy threats when revealed, as one can reasonably reconstruct the image.
@@ 2.2. Visual Localization
In the field of computer vision, the PnP problem (Lepetit et al., 2009) consists of estimating the pose (rotation and translation) of a camera with respect to a world coordinate system. The pose is estimated by exploiting a set of 2D feature locations (e.g. points in an image) and their corresponding 3D points (e.g. a 3D map of the environment). If a camera can move freely in 3D space its pose has six degrees of freedom, three rotational and three translational, about the three Cartesian axis. Applications in robotics and AR/VR often use PnP solvers as part of larger processing pipelines including control (Chaumette and Hutchinson, 2006) and mapping (Schonberger and Frahm, 2016). This work focuses only on pose estimation with the PnP problem.
@@ 2.3. Privacy-Preserving Localization
Engelsma et al. (Engelsma et al., 2022) demonstrate a homomorphically encrypted representation search that could be further adapted to existing visual localization algorithms. This represents a promising frontier, but is currently in the early stages of application and does not scale to such processes in its current form.
The performance envelope of privacy preserving localization proposed in this work is not well suited for time-critical applications like AR/VR, as we later show. There are, however, better suited applications which are less time-sensitive such as Google Visual Positioning System, Facebook Livemaps, and Microsoft Azure Spatial Anchors (Google, 2019; Meta, 2019; Microsoft, 2020). These services offer maps supporting localization for thier users. Privacy preserving localization would allow a user to localize an image or video using a mapping service without the user sharing their image or location for privacy reasons, and without the map provider needing to share their map for commercial reasons.
As another example of where privacy is important in localization, consider a delivery robot which must navigate private facilities to deliver packages. The robot is not trusted to learn the full map of these facilities, nor is it allowed to share images it has taken as they may reveal people’s faces or confidential information from the facility. Privacy preserving localization keeps images on device and the full map of the facilities off device. The robot may run MPC-based localization with two map servers where one is owned by the robot’s manufacturer and the other by the facility who uses their confidential map as input to the protocol. The robot does not learn the full map of the facility, nor do the map servers learn the robot’s location or images.
@@ 2.4. Homomorphic Encryption and Localization
The localization algorithms considered are particularly challenging to adapt to execution via homomorphic encryption (HE). Most problematic, is that localization is a high depth computation with significant data dependencies between steps. For example, to perform one iteration of Levenberg-Marquardt localization, over 1,000 ciphertext-ciphertext divisions and over 7,000 multiplications are required in a high depth circuit, as we later show in Figure 4. The high degree of dependencies between input data does not suit batching based optimizations for HE schemes popular in BGV, BFV and CKKS where multiple plaintexts are encrypted and operated on together (Brakerski et al., 2012; Fan and Vercauteren, 2012; Cheon et al., 2017). This, and the large fractional precision required, suggests localization would best suited for a boolean-based FHE scheme. A popular implementation of one such scheme is tfhe-rs which claims to evaluate a boolean gate in 8.5 ms (Zama, 2022). Given there are 4979 AND and XOR gates per floating point multiplication, for example in EMP-Toolkit’s floating point circuits (Wang et al., 2016), the time required to perform one iteration of Levenberg-Marquardt based localization is estimated to be at least 423 seconds for the multiplications alone. For these reasons, homomorphic encryption is not further considered in favor of an MPC-based approach. While MPC has higher communication complexity than HE, i.e. communication is proportional to the size of the computation vs the size of just the inputs, we later show this cost is practical.
@@ 3. Security Model
This definition gives insight into where previous privacy preserving localization attempts have failed. Line cloud transformations of Speciale et. al. cannot be simulated; the lines contained within the view are not known to the simulator and thus the difference between the real and simulated views can be easily distinguished, as practically proven by Chelani et. al. (Speciale et al., 2019; Chelani et al., 2021). Granted, line clouds were only meant to prevent image reconstruction and do not attempt to hide map feature locations M or the output pose $\mathbf{x}$ (see Table 1). However, just considering input feature locations, an adaptation of the above definition to remove the map M and output pose $\mathbf{x}$ still suggests line cloud transformations are not secure. Furthermore, line cloud transformations do not consider the streaming model; authors note their technique is inversely proportional to the amount of data processed in series. A concrete definition of security is needed to capture repeated invocation via a streaming model to avoid this leakage.
@@ 4. Overview of Privacy Preserving Localization
The second setting considered is where a client holds the features extracted from their image $\mathbf{I}$, but the 3D map $\mathbf{M}$ is held by a third party. The client and third party interact to compute the pose. The client should not learn anything about the map other than what can be inferred from the pose, and the third party should learn (informally) nothing. In the real world, the full map $\mathbf{M}_{\text{full}}$ likely contains more features than the client’s image $\mathbf{I}$. In order to keep $\mathbf{M}_{\text{full}}$ confidential, we assume the existence of an alignment functionality to match the client’s features in $\mathbf{I}$ to the subset of features shared by the map $\mathbf{M}_{full}$, specifically, $\mathbf{M}=align(\mathbf{I},\mathbf{M}_{full})$. We believe the alignment functionality is a reasonably straightforward application of private information retrieval (Chor et al., 1998) and leave concrete instantiation to future work. When alignment is composed with localization, i.e. $x=l(\mathbf{I},\text{align}(\mathbf{I},\mathbf{M}_{\text{full}}))$, the client learns neither the exact map features $\mathbf{M}$, nor map features which did not have a corresponding image feature, i.e. $\{x:x\in\mathbf{M}_{\text{full}}\text{ and }x\notin\mathbf{M}\}$.
In both settings considered, the core challenge to instantiating $l$ in a privacy preserving way is in addressing the iterative nature of localization. Iterative algorithms like localization are not friendly to running under MPC because they are not data-oblivious, meaning control flow depends on secret data i.e. the convergence criteria. To run an iterative algorithm under MPC, an upper bound of iterations must be specified a priori, regardless of convergence at runtime, to ensure the rate of convergence is not leaked. If the rate of convergence were leaked, it reveals sensitive information allowing the devices position to be inferred, especially when localization is run repeatedly like on sequential camera frames (as is common in robotics, AR/VR, and autonomous vehicles). The iteration leakage problem has been studied previously in the context of privacy preserving SAT solvers (Luo et al., 2022; Toft, 2009) and there is no direct solution; an upper bound of iterations must be executed to hide the real number of iterations required for convergence. Localization is particularly problematic as it contains an outer gradient descent iterative algorithm, each iteration of which computes the singular value decomposition (SVD) to (pseudo) invert a matrix which is itself an iterative algorithm. Our Single Iteration Localization approach addresses this, instantiated using an appropriate localization algorithm and MPC protocol given the context.
@@ 5.1. Localization Algorithm
LM is selected for its wide popularity, real-world usage, and generality as it is the default PnP algorithm in OpenCV (Bradski, 2000) and accepts any number of input point correspondences making it useful in combination with other localization related tasks like RANSAC (Fischler and Bolles, 1981). GN is considered as it requires the same set of linear algebra operations as LM, thus is easy to implement, and also requires fewer matrix multiplications at the cost of inverting a larger matrix. While there are many more algorithms, LM and GN are the most popular and require a narrow variety of operations, i.e., matrix arithmetic 2, point projection 3, Euler finite rotation transformation 3, and a matrix (pseudo) inverse via SVD 4. Furthermore, these algorithms can be directly applied, or easily adjusted, to other computer vision problems like relative pose estimation (Bartoli and Sturm, 2004), homography & image alignment (Hartley and Zisserman, 2004), and Structure-from-Motion (SfM) (Schonberger and Frahm, 2016).
@@ 5.2. MPC Library
5.3. Data Representation
Given promising results from using alternative ways to represent data in the field of machine learning, we consider various representations in privacy preserving localization (Gupta et al., 2015; Lesser et al., 2011). Plaintext arithmetic operations on floating point data are accelerated on modern hardware making localization on floating point data fast (around 10 ms for 10k operations) and thus other representations like fixed point are not usually considered. In contrast, when performing floating point operations under the MPC frameworks considered, there are no floating point functional units. Garbled circuit logic gates are evaluated in the same manner no matter how data is represented. For this reason we consider both fixed and floating point data types of various width for privacy preserving localization.
@@ 5.4. Garbled Circuits for Computation Offload
Garbled circuit protocols are often considered to run between two parties, each of which has secret inputs to the function being computed. In the offload setting we consider, instead the offload servers have no inputs to the function and are not allowed to learn the output. Achieving this requires some minor logistical modifications to the garbled circuits protocol setup.
Recall two party garbled circuit protocols have two roles, a generator and an evaluator. The generator creates garbled circuit, which in practice mostly consists of repeatedly evaluating the AES blockcipher. The garbled circuit, i.e. a set of ciphertexts, is then sent to the evaluator who decrypts parts of the received ciphertext, a process that again mostly consists of evaluating AES. The evaluator learns a subset of the ciphertext representing inputs called input wire labels via a primitive named oblivious transfer. The evaluator then uses the input wire labels to decrypt the garbled circuit.
In the setting where all plaintext inputs and outputs are owned by the client, for whom we additionally want to minimize computation and network usage, the simplest way to get the correct input wire labels from generator to evaluator is to have the generator send all labels to the client, who then forwards the appropriate label per wire onwards to the evaluator. We make a slight optimization from the above approach to reduce the bandwidth requirements of the client. The client chooses a random seed and sends it to the generator who uses a pseudorandom function and the seed to create the wire labels for the garbled circuit. Note that the EMP toolkit framework already uses a seeded PRF to generate the garbled circuit for performance reasons (/dev/random is slow); we have instead allowed the client to choose this seed which is not an issue since the client is the only party who supplies the secret input. After, the generator creates the garbled circuit, they send it to the evaluator excluding the input wire labels. The client, knowing the circuit seed, can generate the input wire labels without interaction and send them to the evaluator. The circuit is evaluated as usual and the output labels are sent back to the client who can decode their plaintext meaning. This saves the client from needing to receive all wire labels from the generator. Instead, the client may generate them directly and send to the evaluator. Since the same seed is used to generate input wire labels and the rest of the garbled circuit, correctness holds. Note in the case of EMP and the half-gates protocol, the client must also know the global circuit delta (Kolesnikov and Schneider, 2008) as part of the generation process. Intuitively, it is okay for the client to learn the seed and delta as the client is the only party who supplies secret inputs and is allowed to learn the output. While this is appropriate for the offload setting, in the setting where the input image and map are known to different parties, this optimization is not appropriate.